4.(2 pts) A computer lab consultant has to check the functionality of the printers in a certain building. Suppose that 7% of printers have defects, and defects occur independently of each other. Find the probability of exactly 2 defective printers in the building that has thirty printers.
Answer #1

Here, n = 30, p = 0.07, (1 - p) = 0.93 and x = 2
As per binomial distribution formula P(X = x) = nCx * p^x * (1 - p)^(n - x)

We need to calculate P(X = 2)
P(X = 2) = 30C2 * 0.07^2 * 0.93^28
P(X = 2) = 0.2794